# A proof-of-principle experiment on the spontaneous symmetry breaking machine and numerical estimation of its performance on the benchmark problem

### v1

@@ IV.2 Competition between dynamics-derived attractions and PSI
Figure 10 shows the numerical simulation results for obtained using Eq.(8), illustrating an example where the PSI was set quite large. Note that the behavior of the IE with respect to the step (see Fig.10(b)) was calculated using results where was classified as 1-state and as 0-state. As can be seen from these figures, the competition between the attraction derived from dynamics and PSI causes the pseudo-spin to fail to sufficiently attract toward the 0-state/1-state in regions where the PSI value is large. Unlike Ising model-based simulators (such as CIM or SBM), SSBM guaranties behavior based on the interpretation of exchange interaction energy even in such situations. Therefore, it is possible to promote stabilization as measured by the IE, and the behavior shown in Fig.10 also demonstrates this fact. However, when considering the practical operation of SSBM, it is anticipated that each pseudo-spin will become an obstacle in the process of determining whether it converged to the 0-state or 1-state. Furthermore, the fact that the stable state where the pseudo-spin is extremely far from the 0-state & 1-state is displaced from the Ising energy potential surface raises concerns about the degradation of performance as a simulator for the COP, which requires finding stable states as IE. Therefore, we investigated ways to enhance the attraction to 0-state&1-state arising from dynamics and evolve the system into one capable of achieving proper state separation of pseudo-spins (namely, ensuring that the pseudo-spin converges sufficiently to the 0-state&1-state) against competing PSI. As a result, we arrived at an evolved version of SSBM expressed by the following equations.
@@
Figure 12 shows the results of the numerical simulation for the case where the NNA was controlled. In the case where the number of nested actions was increased by one (fixed), it was confirmed, as initially intended, that the separation of the 0-state & 1-states of the pseudo-spin could be effectively achieved (Fig.12(a)) and that stabilization in terms of IE was improved (Fig.12(b)). Furthermore, in this case, it was observed that the variance in the IE in the final state (step300) was relatively large. This is thought to occur because increasing the NNA rapidly advances both the convergence of the pseudo-spins toward the attractor and the fixation of the state. On the other hand, in the case of dynamic control where the NNA is not increased until step=100 and increases once thereafter, not only the state separation of pseudo-spins in the final state (Fig.12(c)) and the improvement in stabilization (Fig.12(d)), but also the suppression of variation in IE was confirmed (Fig.12(d)). This can be considered evidence that, during the initial 100 steps, there is not only the disadvantage that pseudo-spin state separation does not progress, but also the advantage that transitions to more stable pseudo-spin configurations are occurring gradually yet reliably. Furthermore, it is thought that this advantage is realized by PSI-specific effects not present in exchange interactions, suggesting that effective control of NNA could enable more advanced stabilization control (solution exploration).
@@ IV.3 Numerical Simulation of SSBM Applied to Problem
Figure 13 shows the change in cut values (CV) with a step progression for 1000 numerical simulations. Since many reports dealing with large-scale problems use CVs rather than IEs, we followed this convention. Additionally, as calculation conditions, was adjusted within the range of , while the NNA was incrementally increased from to , and Gaussian random numbers [25] differing for each simulation were used as fluctuations. Specifically, in the case of Fig.13(a) (case-1), for and , the number of steps was set to and , respectively. Then in the case of Fig.13(b) (case-2), only the number of steps for was increased from to . And, Figs.14(a)-(c) show the histograms of the CVs for 1250, 1400, and 3750 steps in case-1, respectively, while Figs. 14(d)-(f) show those for 2000, 2150, and 4500 steps in case-2, respectively. These simulation results demonstrate that SSBM possesses distinct characteristics that distinguish it from the other COP solvers. For example, in the case of quantum computers (quantum annealers), the tunneling effect can only be expected with a certain probability within a realistic computational time (finite time). Consequently, statistical variation in the obtained solution (CVs) is unavoidable. Furthermore, statistical variation in the CVs obtained has also been confirmed in reported cases involving other Ising machines, including algorithmic types [26, 17, 19, 20]. On the other hand, SSBM exhibited an extremely unique and interesting behavior in which a single stable state is created (Figure 13(b)), and the CV observed there indicated that SSBM has top-level characteristics (99.7% of the best value currently known). And, this single stable state was confirmed to be paired with a state where the 1-state and 0-state are exactly reversed, and it was also confirmed that both states could be obtained.
@@
We consider this distinctive characteristic of SSBM to be significant in two respects, one of which is as a key to deepening our understanding of the phenomenon itself (its value as a science). We believe that the SSB phenomenon (process) employed in SSBM, and furthermore the newly introduced PSI in SSBM, create a significant difference from the Ising machines. For example, in Ising machines, where continuous variables correspond to spins, gradually varying parameters allow the system to progress toward a stable state while remaining close to equilibrium [17, 18, 19]. Broadly speaking, this operation separates all the continuous variables into two groups, up and down, effectively maintaining conditions that keep the system near the Ising energy potential surface. This operation is necessary to induce the transition to a stable state in systems where exchange interactions are introduced directly.
@@
On the other hand, in SSBM, the PSI can function effectively regardless of the relative positional relationship of continuous variables (even when they are not separated into two distinct states corresponding to up and down). And, the attraction toward the 1-state or 0-state inherent in the dynamics themselves competes with PSI, determining the stable state as a pseudo-spin. Consequently, for example, at , it exhibits extremely gradual changes or stagnation while assuming a scattered state over a wide range between the 1-state and 0-state. Moreover, the gradual changes observed here suggest a transition toward a stable state in a particular sense, even though they do little to further advance stabilization from the perspective of the IE.
@@
For example, while there is little difference in the CVs at the final step for between case-1 and case-2 (cf. Fig.14(a) and Fig.14(d)), a significant difference becomes apparent when comparing the 1400-th step of case-1 (Fig.14(b)) with the 2150-th step of case-2 (Fig.14(e)) (150 steps after switching the NNA to ). In the final step, it was confirmed that the stable state obtained in case-1 retains statistical fluctuations (Fig.14(c)), while convergence to a single stable state was observed in case-2 (Fig.14(f)). Furthermore, the following points were also confirmed from the simulation results.
@@
Result-3: The behavior of CV (IE) (Fig.13(a)) reveals that there are streams converging to several specific stable states and that the transition is not merely monotonically toward states with larger CVs (states with lower IE).
@@
Prediction-3:
(Consideration that the CV histogram in the 1450-th step for n=6 in case-2 is identical to Fig.14(a))
Although no significant change in CV (IE) is observed during the transition from Fig.14(a) to Fig.14(d), it is considered that the state change shifting to stream-one is completed during this period. This state change is driven by the entropy effect arising from the relatively large number of states constituting stream-one within this energy range.
Verification from a different perspective to reinforce the above predictions remains a future task.
Another important consideration is the advantages and disadvantages of its application as a COP solver. For example, when considering application scenarios for the COP solvers (quantum anealer, CIM, SBM and STATICA etc.) where the obtained solutions exhibit statistical variability, it is necessary to determine the number of computational samples and further consider the need for subsequent comparison and selection processes for the obtained sample solutions. And, it will also be important to determine the extent to which the device configuration can support parallel computing. On the other hand, since SSBM can converge to a single solution in the final step, a single computational sample is sufficient, and subsequent comparison and selection processes are unnecessary. This can be considered a major advantage in terms of compressing computational resources. However, the fact that the solution obtained in SSBM does not currently reach the best cut value known at this time is indeed a disadvantage. However, since this single solution has top-level characteristics (99.7% of the currently known best value), it is considered to offer sufficient potential for practical application and could serve as a viable option. Furthermore, it is important to note that the scientific understanding of the phenomena leading to this single solution and the examination of its impact on applications remain major challenges for the future.
@@ V Conclusion
We have successfully demonstrated experimentally the operation of SSBM and confirmed its promise as a new physically implemented type COP solver. This result also strongly supports the existence of a duality (correspondence) between the IMLS states that appear in SSBM and the stable states of the IMS. In addition, the characteristics were evaluated by numerical simulation on a benchmark problem (), assuming application to large-scale problems. The results of 1000 simulations with different fluctuations confirm that SSBM is able to search for states corresponding to 99.7% of the best cut value currently known, without statistical variability. This is due to the inherent properties of SSBM, rooted in its principles and not found in other solvers, and is expected to be a significant advantage in applications.

### v2

@@ IV.2 Competition between dynamics-derived attractions and pSI
Figure 10 shows the numerical simulation results for obtained using Eq.(8), illustrating an example where the pSI was set quite large. Note that the behavior of the Ising energy with respect to the step (see Fig.10(b)) was calculated using results where was classified as 1-state and as 0-state. As can be seen from these figures, the competition between the attraction derived from dynamics and pSI causes the pseudo-spin to fail to sufficiently attract toward the 0-state/1-state in regions where the pSI value is large. Unlike Ising model-based simulators (such as coherent Ising machine or simulated bifurcation machine), SSBM guarantees behavior based on the interpretation of exchange interaction energy even in such situations. Therefore, it is possible to promote stabilization as measured by the Ising energy, and the behavior shown in Fig.10 also demonstrates this fact. However, when considering the practical operation of SSBM, it is anticipated that each pseudo-spin will become an obstacle in the process of determining whether it converged to the 0-state or 1-state. Furthermore, the fact that the stable state where the pseudo-spin is extremely far from the 0-state & 1-state is displaced from the Ising energy potential surface raises concerns about the degradation of performance as a simulator for the COP, which requires finding stable states as Ising energy. Therefore, we investigated ways to enhance the attraction to 0-state&1-state arising from dynamics and evolve the system into one capable of achieving proper state separation of pseudo-spins (namely, ensuring that the pseudo-spin converges sufficiently to the 0-state&1-state) against competing pSI. (Although studies on improving convergence to up-state and down-state have been reported for Ising machines utilizing bifurcation phenomena [26, 31], the SSBM, which employs dynamics-derived attraction, required an entirely different approach.) As a result, we arrived at an evolved version of SSBM expressed by the following equations.
@@
Figure 12 shows the results of the numerical simulation for the case where the NNA was controlled. In the case where the number of nested actions was increased by one (fixed), it was confirmed, as initially intended, that the separation of the 0-state & 1-states of the pseudo-spin could be effectively achieved (Fig.12(a)) and that stabilization in terms of Ising energy was improved (Fig.12(b)). Furthermore, in this case, it was observed that the variance in the Ising energy in the final state (step300) was relatively large. This is thought to occur because increasing the NNA rapidly advances both the convergence of the pseudo-spins toward the attractor and the fixation of the state. On the other hand, in the case of dynamic control where the NNA is not increased until step=100 and increases once thereafter, not only the state separation of pseudo-spins in the final state (Fig.12(c)) and the improvement in stabilization (Fig.12(d)), but also the suppression of variation in Ising energy was confirmed (Fig.12(d)). This can be considered evidence that, during the initial 100 steps, there is not only the disadvantage that pseudo-spin state separation does not progress, but also the advantage that transitions to more stable pseudo-spin configurations are occurring gradually yet reliably. Furthermore, it is thought that this advantage is realized by pSI-specific effects not present in exchange interactions, suggesting that effective control of NNA could enable more advanced stabilization control (solution exploration).
@@ IV.3 Numerical Simulation of SSBM Applied to Problem
Figure 13 shows the change in cut values with a step progression for 1000 numerical simulations. Since many reports dealing with large-scale problems use cut values rather than Ising energys, we followed this convention. Additionally, as calculation conditions, the NNA was incrementally increased from to , and the value of was adjusted to gradually decrease within a narrow range of for each NNA value, and Gaussian random numbers [32] (with a standard deviation of ) differing for each simulation were used as fluctuations. Specifically, in the case of Fig.13(a) (case-1), for and , the number of steps was set to and , respectively. Then in the case of Fig.13(b) (case-2), only the number of steps for was increased from to . And, Figs.14(a)-(c) show the histograms of the cut values for 1250, 1400, and 3750 steps in case-1, respectively, while Figs. 14(d)-(f) show those for 2000, 2150, and 4500 steps in case-2, respectively. The simulation parameters used here were identified through repeated trials. Future challenges include developing methods to find parameters more efficiently using AI technology or other techniques, as well as examining their applicability to a broader range of problem types. These simulation results demonstrate that SSBM possesses distinct characteristics that distinguish it from the other COP solvers. For example, in the case of quantum computers (quantum annealers), the tunneling effect can only be expected with a certain probability within a realistic computational time (finite time). Consequently, statistical variation in the obtained solution (cut values) is unavoidable. Furthermore, statistical variation in the cut values obtained has also been confirmed in reported cases involving other Ising machines, including algorithmic types [33, 17, 19, 20]. On the other hand, SSBM exhibited an extremely unique and interesting behavior in which a single stable state is created (Figure 13(b)), and the cut value observed there indicated that SSBM has top-level characteristics (99.7% of the best value currently known [19]). And, this single stable state was confirmed to be paired with a state where the 1-state and 0-state are exactly reversed, and it was also confirmed that both states could be obtained.
@@
We consider this distinctive characteristic of SSBM to be significant in two respects, one of which is as a key to deepening our understanding of the phenomenon itself (its scientific value). We believe that the SSB phenomenon (process) employed in SSBM, and furthermore the newly introduced pSI in SSBM, create a significant difference from the Ising machines. For example, in Ising machines, where continuous variables correspond to spins, gradually varying parameters allow the system to progress toward a stable state while remaining close to equilibrium [17, 18, 19]. Broadly speaking, this operation separates all the continuous variables into two groups, up and down, effectively maintaining conditions that keep the system near the Ising energy potential surface. This operation is necessary to induce the transition to a stable state in systems where exchange interactions are introduced directly.
@@
On the other hand, in SSBM, the pSI can function effectively regardless of the relative positional relationship of continuous variables (even when they are not separated into two distinct states corresponding to up and down). And, the attraction toward the 1-state or 0-state inherent in the dynamics themselves competes with pSI, determining the stable state as a pseudo-spin. Consequently, for example, at , it exhibits extremely gradual changes or stagnation while assuming a scattered state over a wide range between the 1-state and 0-state. Moreover, the gradual changes observed here suggest a transition toward a stable state in a particular sense, even though they do little to further advance stabilization from the perspective of the Ising energy.
@@
For example, while there is little difference in the cut values at the final step for between case-1 and case-2 (cf. Fig.14(a) and Fig.14(d)), a significant difference becomes apparent when comparing the 1400-th step of case-1 (Fig.14(b)) with the 2150-th step of case-2 (Fig.14(e)) (150 steps after switching the NNA to ). In the final step, it was confirmed that the stable state obtained in case-1 retains statistical fluctuations (Fig.14(c)), while convergence to a single stable state was observed in case-2 (Fig.14(f)). Furthermore, the following points were also confirmed from the simulation results.
@@
Result-3: The behavior of cut value (Ising energy) (Fig.13(a)) reveals that there are streams converging to several specific stable states and that the transition is not merely monotonically toward states with larger cut values (states with lower Ising energy).
@@
Prediction-3: (Consideration that the cut value histogram in the 1450-th step for n=6 in case-2 is identical to Fig.14(a)) Although no significant change in cut value (Ising energy) is observed during the transition from Fig.14(a) to Fig.14(d), it is considered that the state change shifting to stream-one is completed during this period. This state change is driven by the entropy effect arising from the relatively large number of states constituting stream-one within this energy range.
In an analog variable-type Ising machine that directly inherits Ising interactions using branching phenomena, operations are typically performed to slowly increase the gain and feedback strength so that the steady-state approximation or adiabatic approximation holds. This maintains a state separated into two states, upper-side and lower-side, while approaching the up-spin or down-spin state. In such a process, the system does not fall into the state shown in Figure 10(a). However, if forced into a similar configuration, the interaction itself exhibits behavior resembling significant noise. The system then undergoes a behavior that resolves this state and settles into the separated upper-side and lower-side states. In contrast, SSBM, by introducing a pseudo-spin interaction, guarantees behavior consistent with the interpretation of exchange interaction energy even when the system is not separated into two states (e.g., as in Figure 10(a)), maintaining this state. Noting that the only difference in conditions is the number of steps for n=6, the above results suggest that the unique characteristic of SSBM in the region not separated into these two states effectively led to pattern convergence by ensuring a sufficient number of steps. Verification from a different perspective to reinforce the above predictions remains a future task.
Another important consideration is the advantages and disadvantages of its application as a COP solver. For example, when considering application scenarios for the COP solvers (quantum anealer, coherent Ising machine, simulated bifurcation machine and STATICA etc.) where the obtained solutions exhibit statistical variability, it is necessary to determine the number of computational samples and further consider the need for subsequent comparison and selection processes for the obtained sample solutions. And, it will also be important to determine the extent to which the device configuration can support parallel computing. On the other hand, since SSBM holds promise for effectively converging patterns that can be obtained under certain conditions, we anticipate that gaining broader insights into the relationship between simulation conditions and convergence effects could lead to discovering efficient utilization methods that conserve computational resources. While it is true that the solutions obtained via SSBM have not yet reached the current best cut value, we consider it a promising option since the solutions achieved represent top-level performance (99.7% of the currently known best value). Furthermore, it is important to note that the scientific understanding of the phenomena leading to this single solution and the examination of its impact on applications remain major challenges for the future.
@@ V Conclusion
We have successfully demonstrated experimentally the operation of SSBM and confirmed its promise as a new physically implemented type COP solver. This result also strongly supports the existence of a duality (correspondence) between the Ising-model-like system states that appear in SSBM and the stable states of the Ising model system. In addition, the characteristics were evaluated by numerical simulation on a benchmark problem (), assuming application to large-scale problems. The results of 1000 simulations with different fluctuations confirm that SSBM is able to search for states corresponding to 99.7% of the best cut value currently known, without statistical variability. This is a unique characteristic rooted in SSBM’s principles, not found in other solvers, and has the potential to become a distinct advantage in the future. However, in configurations that realize all individual pSIs presented here via optical interference, as the number of pSIs increases, the optical power of each individual pSI must be reduced, making the SN ratio of the optical pulses the root cause of a severe scalability problem. Similarly, for coherent Ising machine, a solver also using optical pulses, no experimental reports exist of configurations achieving Ising interactions solely through optical interference, even for problems where all spins are coupled. This is because coherent Ising machine faces challenges similar to those of SSBM. This scalability problem is expected to be solvable by following the example of coherent Ising machine, which succeeded in experimental application to the fully-connected problem with N=100000 [34]. This scalability issue is expected to be resolved by following the example of coherent Ising machine, which successfully applied to the N=100,000 fully connected problem. There, the problem is addressed using a method termed a measurement and feedback scheme: optically encoded signals corresponding to spins are converted to electrical signals, then interactions are computed digitally using an FPGA, and finally the computed interactions are feedback as optical signals via electro-optical conversion.